# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### v3

@@ 3.3 Data augmentation
We utilized frame interpolation model to augment bradycardia and tachycardia data, and Fig. 2 presents the data augmentation process in temporal aspects. Normal heart rates generally range from 60 BPM to 100 BPM [Guyton, 2006], and bradycardia and tachycardia refer to heart rates lower and higher than the normal heart rate. In the training process, an ROI sequence of the length $T$ corresponding to the heart rate $r$ beats per minute (BPM) is augmented to generate bradycardia and tachycardia data which correspond to the heart rates $\frac{r}{2}$ BPM and $2r$ BPM, respectively. To augment the bradycardia data, $\frac{T}{2}$ frames of the original ROI sequence is interpolated into the length of $T$ frames by utilizing a pretrained FILM-Net [Reda et al., 2022]. The duration of a cardiac cycle increases as the interpolation rate increases, resulting in decreased heart rate. In contrast, $2T$ frames of the original ROI sequence is downsampled with the factor of $2$ to augment tachycardia data. Previous data augmentation methods are based on frame sampling [Yu et al., 2020, Lokendra & Puneet, 2022], color jittering [Yue et al., 2023], and ROI masking [Chen et al., 2023]. Different to these previous approaches, data augmentation based on a frame interpolation model has the benefit of being able to directly control the heart rate of the augmented data.
Figure 2: Augmentation of bradycardia and tachycardia data.
3.4 Dual remote photoplethysmography network (DRP-Net)
In this paper, we propose DRP-Net to estimate phase-shifted rPPG signals from image sequences, and Fig. 3 presents the architecture of the DRP-Net. DRP-Net takes a sequence of facial images $\textbf{X}\in\mathbb{R}^{3\times T\times H\times W}$, where $H$ and $W$ are height and width, and $T$ is the number of frames in a window. In experiments, $H$ and $W$ are set to 128, and $T$ is set to 150. DRP-Net is a 3D CNN model, consisting of a feature extractor, spatial and temporal attention modules, and rPPG prediction heads. The feature extractor consists of atrous convolution layers to analyze spatiotemporal features over a large receptive field. In Fig. 3, the kernel size and dilation ratio for each convolution layer are denoted by $k$ and $d$; in a 3D convolution layer, $k$ and $d$ are tuples with the form of [$T$, $H$, $W$], and it is represented as a scalar if three values are same. The feature extractor consists of seven atrous convolution layers and three max pooling layers, and it produces a spatiotemporal feature map $\textbf{m}\in\mathbb{R}^{64\times T\times 4\times 4}$.
@@ 3.4 Dual remote photoplethysmography network (DRP-Net)
The intermediate feature map $\textbf{m}_{inter}\in\mathbb{R}^{64\times T\times 32\times 32}$ of the feature extractor is utilized to compute spatial and temporal attention in parallel. The spatial attention module consists of 3D convolution, max pooling, average pooling, and pointwise convolution layers. The sigmoid function $\sigma$ is applied on the attention score map $\textbf{v}_{s}\in\mathbb{R}^{1\times 1\times 4\times 4}$ to produce the spatial attention $\boldsymbol{\alpha}_{s}$. Similarly, in the temporal attention module, we compute a temporal score vector $\textbf{v}_{t}\in\mathbb{R}^{1\times T\times 1\times 1}$, and the sigmoid function is applied to obtain the temporal attention vector $\boldsymbol{\alpha}_{t}$. The spatial and temporal attention modules refine the spatiotemporal feature map m as follows.
@@
3.5 Bounded blood pressure network (BBP-Net)
BBP-Net is designed to estimate SBP and DBP values from facial and acral rPPG signals, and its structure is presented in Fig. 4. BBP-Net takes a stack of physiological signals consisting of phase-shifted rPPG signals, velocity plethysmography (VPG) and acceleration plethysmography (APG) signals. VPG and APG are the first and second derivatives of the facial and acral rPPG signals. BBP-Net consists of pointwise convolution layer, MSF blocks, BP prediction heads, and scaled sigmoid layers.
The MSF block integrates physiological information from various receptive fields. In Fig. 4, $C_{in}$, $C_{out}$, and $C_{mid}$ denote the channels of input, output, and middle layers in MSF blocks and BP prediction heads; $T_{in}$ denotes the temporal length of an input signal. The MSF block performs depth-wise separable (DWS) convolution with the kernel sizes of 3 and 5 in parallel, and it employs the Hardswish activation function [Howard et al., 2019]. After the global average pooling, the global feature vector with channel $C_{mid}$ passes linear layers and a softmax layer to compute weight vectors for the outputs of two DWS convolution layers. MSF blocks contain a residual connection to reduce the problem of vanishing gradients.
The BP prediction head includes a bottleneck attention module [Park et al., 2018] and residual connections to extract temporal features. Moreover, we employed a scaled sigmoid function to constrain the estimated blood pressure into a predefined range as follows.
@@ 3.6 Loss function
$S_{p}(\textbf{y})=\{t\mid(\textbf{y}(t)-\textbf{y}(t-1))(\textbf{y}(t+1)-\textbf{y}(t))<0,~{}\textbf{y}(t)>\textbf{y}(t-1)\},$ (6)
@@
$S_{v}(\textbf{y})=\{t\mid(\textbf{y}(t)-\textbf{y}(t-1))(\textbf{y}(t+1)-\textbf{y}(t))<0,~{}\textbf{y}(t)<\textbf{y}(t-1)\}.$ (8)
@@
For the training of BBP-Net, we define $L_{BP}$ and $L_{ABP}$. $L_{BP}$ employs Huber loss [Collins, 1976] to calculate the loss of predicted SBP and DBP, and it is defined as the following equation.
@@
where $\hat{y}_{max}$ and $\hat{y}_{min}$ are the maximum and minimum values of $\hat{\textbf{y}}_{a}$. $L_{ABP}$ is defined as the L2 distance between the scaled acral rPPG signal $\hat{\textbf{y}}_{s}$ and its corresponding ABP signal Y as follows.
$L_{ABP}=\lVert\hat{\textbf{y}}_{s}-\textbf{Y}\rVert_{2}.$ (17)
@@ 4 Experimental results
Experiments were conducted on a hardware environment including Intel Core i9-10940X CPU, 64 GB DDR4 RAM, and NVIDIA Geforce RTX 3090 Ti. Python and Pytorch were utilized to implement the proposed algorithm, and code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG. In experiments, the temporal window was set to 150 samples which corresponds to 6 seconds. The learning rates of the Adam optimizer were set to 0.001 and 0.0001 for the MMSE-HR and V4V database, and the batch size was set to 8. To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE), root mean squared error (RMSE), and Pearson correlation coefficient r as evaluation measures.
4.1 Datasets
@@
We conducted experiments using the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) [Zhang et al., 2016b] and V4V (Vision for Vitals) database [Revanur et al., 2021]. The datasets are sub-datasets derived from the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous ABP signals. As explained in Section 3.2, pseudo PPG signals, heart rate, SBP, and DBP were obtained from ABP signals. In the acquisition of the datasets, subjects were conducted various tasks to arouse amusement, pain, anger, and disgust. RGB video was recorded at the resolution of 1040$\times$1392 with the frame rate of 25 FPS, and ABP signals were collected at the sampling rate of 1000 Hz. The MMSE-HR database includes 102 video sequences from 17 male and 23 female subjects, and the average length of sequences is 30 seconds. The V4V database includes 1,358 data sequences from 179 subjects, and they are split into 724 for training, 276 for validation, and 358 for test samples. The average length of the sequences in the V4V database is 40 seconds.
4.2 Experimental results on MMSE-HR database
@@
The performance of the DRP-Net for estimating heart rate is compared with previous methods on the MMSE-HR dataset. Following the previous work [Yu et al., 2023], experiments on the MMSE-HR dataset were conducted by using 5-fold cross-validation method for heart rate estimation. The folds were split independently between subjects to demonstrate generalizability of the proposed method. Table 1 presents the averaged performance over 5 folds for estimating heart rate from facial and acral rPPG signals. The proposed method achieved MAE of 1.78, RMSE of 4.27, and r of 0.95, respectively, outperforming previous deep learning models with a significant margin. Previous methods proposed by Nowara et al. [2021], Jaiswal & Meenpal [2022], Ouzar et al. [2023] infer rPPG signals based on small window sizes with low latency. However, insufficient temporal information leads to increase errors in estimating rPPG signals, resulting in higher heart rate estimation errors. Video transformer-based models proposed by Yu et al. [2022] and Yu et al. [2023] demonstrate significant improvements in cross-dataset tests. However, transformers require high-performance computing resources due to their huge amount of parameters and computational demands. As shown in Table 1, our proposed model achieved better performance compared to the previous methods by using similar length of video sequences. In particular, the estimated facial rPPG signals showed the lowest MAE for estimating heart rate in the MMSE-HR dataset. These results imply that forcing the phase of the estimated rPPG signal to be same as the acral PPG signal may result in decreasing the performance of heart rate estimation.
Table 1: Heart rate estimation results on the MMSE-HR database.
Method Window (sec) MAE (BPM) RMSE r
POS [De Haan & Jeanne, 2013] - 5.77 - 0.82
DeepPhys [Chen & McDuff, 2018] - 4.72 8.68 0.82
Benefit of distraction [Nowara et al., 2021] 2 2.27 4.90 0.94
EfficientPhys-C [Liu et al., 2021] - 3.48 7.21 0.86
CAN with synthetic data [McDuff et al., 2022] 30 2.26 3.70 -
PhysFormer [Yu et al., 2022] 6.4 2.84 5.36 0.92
Spatiotemporal feature [Jaiswal & Meenpal, 2022] 5 6.40 6.82 0.95
X-iPPGNet [Ouzar et al., 2023] 2 4.10 5.32 0.85
PhysFormer++ [Yu et al., 2023] 6.4 2.71 5.15 0.93
CIN-rPPG [Li et al., 2023] 12 1.93 4.43 0.94
Dual-TL [Qian et al., 2024] 12 2.25 4.27 0.93
Ours (facial) 6 1.78 4.27 0.95
Ours (acral) 6 1.91 4.74 0.93
@@
Table 2 presents the results of estimating blood pressure on the MMSE-HR database. The proposed BBP-Net for blood pressure estimation outperformed previous methods, achieving MAE and RMSE of 10.19 and 13.01 for SBP and 7.09 and 8.86 for DBP, respectively. The algorithms proposed by Rong & Li [2021] and Schrumpf et al. [2021] extracted rPPG signals using a non-parametric approach and estimated blood pressure based on handcrafted features and deep learning models. These previous studies are similar to our approach in the aspect of extracting rPPG signals in an intermediate step. However, the main difference of our proposed method is to analyze phase-shifted rPPG signals extracted from DRP-Net. Chen et al. [2023] proposed a blood pressure estimation model that utilizes two-dimensional spatiotemporal maps obtained from facial videos. However, this previous method cannot analyze pulse transit time of rPPG signals and shows insufficient performance for estimating SBP and DBP. Experimental results in Table 2 demonstrate that analyzing temporal discrepancy in phase-shifted rPPG signals is meaningful for improving the precision of blood pressure estimation. Fig. 6 presents the Bland-Altman plots of predicted and reference blood pressure for the MMSE-HR database. In Fig. 6, solid line and dotted lines represent mean error and 95% limits of agreement (LOA), respectively.
Table 2: Blood pressure estimation results on the MMSE-HR database.
Method Window (sec) SBP DBP
@@
NCBP [Rong & Li, 2021] - 17.52 22.43 12.13 15.23
NIBPP [Schrumpf et al., 2021] 7 13.60 - 10.30 -
BPE-Net [Chen et al., 2023] 6 12.35 16.55 9.54 12.22
@@
4.3 Experimental results on V4V database
This section presents experimental results for estimating heart rate on the V4V database. In Table 3, the performance of our proposed model is compared with previous methods. Facial rPPG signals estimated from the DRP-Net outperformed others in terms of MAE and r, with the values of 3.83 and 0.75, respectively. Similar to the results on the MMSE-HR dataset, our acral rPPG signals show slightly higher error rates compared to the results of the facial rPPG signals. The performance comparison between PhysNet [Yu et al., 2019] and our DRP-Net indicates that utilizing wider receptive fields through atrous convolution is more advantageous for learning spatiotemporal features than the encoder-decoder structure based on three dimensional CNN. Moreover, compared to APNET [Kim et al., 2022], our facial rPPG signals showed lower error rates in terms of MAE and r.
Table 3: Heart rate estimation results on the V4V database.
Method Window (sec) MAE (BPM) RMSE r
DeepPhys [Chen & McDuff, 2018] 30 10.20 13.25 0.45
PhysNet [Yu et al., 2019] - 13.15 19.23 0.75
APNET [Kim et al., 2022] - 4.89 7.68 0.74
Ours (facial) 6 3.83 9.59 0.75
Ours (acral) 6 4.13 10.14 0.73
Table 4: Blood pressure estimation results on the V4V database.
Method Window (sec) SBP DBP

### v4

@@ III-C Data augmentation
We utilized frame interpolation model to augment bradycardia and tachycardia data, and Fig. 2 presents the data augmentation process in temporal aspects. Normal heart rates generally range from 60 BPM to 100 BPM [39], and bradycardia and tachycardia refer to heart rates lower and higher than the normal heart rate. The left histogram in Fig. 2 shows the heart rate distribution of the V4V trainset. The limited heart rate data distribution can constrain the model’s heart rate estimation range, triggering the need for data augmentation. In the training process, an ROI sequence of the length $T$ corresponding to the heart rate $r$ beats per minute (BPM) is augmented to generate bradycardia and tachycardia data which correspond to the heart rates $\frac{r}{2}$ BPM and $2r$ BPM, respectively. To augment the bradycardia data, $\frac{T}{2}$ frames of the original ROI sequence is interpolated into the length of $T$ frames by utilizing a pretrained FILM-Net [40]. The duration of a cardiac cycle increases as the interpolation rate increases, resulting in decreased heart rate. In contrast, $2T$ frames of the original ROI sequence is downsampled with the factor of $2$ to augment tachycardia data. Previous data augmentation methods are based on frame sampling [41, 42], color jittering [43], and ROI masking [36]. Different to these previous approaches, data augmentation based on a frame interpolation model has the benefit of being able to directly control the heart rate of the augmented data.
@@
III-D Dual remote photoplethysmography network (DRP-Net)
In this paper, we propose DRP-Net to estimate phase-shifted rPPG signals from image sequences, and Fig. 3 presents the architecture of the DRP-Net. We designed DRP-Net as a multi-task learning model to extract facial and acral rPPG signals that have similar trends but different phases. DRP-Net takes a sequence of facial images $\textbf{X}\in\mathbb{R}^{3\times T\times H\times W}$, where $H$ and $W$ are height and width, and $T$ is the number of frames in a window. In experiments, $H$ and $W$ are set to 128, and $T$ is set to 150. DRP-Net is a 3D CNN model, consisting of a feature extractor, spatial and temporal attention modules, and rPPG prediction heads. The feature extractor consists of atrous convolution layers to analyze spatiotemporal features over a large receptive field. In Fig. 3, the kernel size and dilation ratio for each convolution layer are denoted by $k$ and $d$; in a 3D convolution layer, $k$ and $d$ are tuples with the form of [$T$, $H$, $W$], and it is represented as a scalar if three values are the same. The feature extractor consists of seven atrous convolution layers and three max pooling layers, and it produces a spatiotemporal feature map $\textbf{m}\in\mathbb{R}^{64\times T\times 4\times 4}$.
The intermediate feature map $\textbf{m}_{inter}\in\mathbb{R}^{64\times T\times 32\times 32}$ of the feature extractor is utilized to compute spatial and temporal attention in parallel. The objective of the spatial and temporal attention modules is to emphasize important regions of the face and the temporal peak locations within the facial video. The spatial attention module consists of 3D convolution, max pooling, average pooling, and pointwise convolution layers. The sigmoid function $\sigma$ is applied on the attention score map $\textbf{v}_{s}\in\mathbb{R}^{1\times 1\times 4\times 4}$ to produce the spatial attention $\boldsymbol{\alpha}_{s}$. Similarly, in the temporal attention module, we compute a temporal score vector $\textbf{v}_{t}\in\mathbb{R}^{1\times T\times 1\times 1}$, and the sigmoid function is applied to obtain the temporal attention vector $\boldsymbol{\alpha}_{t}$. The spatial and temporal attention modules refine the spatiotemporal feature map m as follows.
@@ III-D Dual remote photoplethysmography network (DRP-Net)
III-E Bounded blood pressure network (BBP-Net)
BBP-Net is designed to estimate SBP and DBP values from facial and acral rPPG signals, and its structure is presented in Fig. 4. BBP-Net takes a stack of physiological signals consisting of phase-shifted rPPG signals, velocity plethysmography (VPG) and acceleration plethysmography (APG) signals. VPG and APG are the first and second derivatives of the facial and acral rPPG signals. To extract local temporal features such as phase shifts and signal waveforms, BBP-Net is stacked with convolution-based modules, consisting of pointwise convolution layers, MSF blocks, BP prediction heads, and scaled sigmoid layers.
The MSF block integrates physiological information from various receptive fields. In Fig. 4, $C_{in}$, $C_{out}$, and $C_{mid}$ denote the channels of input, output, and middle layers in MSF blocks and BP prediction heads; $T_{in}$ denotes the temporal length of an input signal. The MSF block performs depth-wise separable (DWS) convolution with the kernel sizes of 3 and 5 in parallel. Since phase-shifted rPPG signals contain negative values, we employed the Hard Swish activation function [45] to handle negative values while preserving the smooth characteristics of the signal data. After the global average pooling, the global feature vector with channel $C_{mid}$ passes linear layers and a softmax layer to compute weight vectors for the outputs of two DWS convolution layers. MSF blocks contain a residual connection to reduce the problem of vanishing gradients.
The BP prediction head includes a bottleneck attention module [44] and residual connections to extract temporal features. Moreover, we employed a scaled sigmoid function to constrain the estimated blood pressure into a predefined range as follows.
@@ III-F Loss function
$S_{p}(\textbf{y})=\{t\mid(\textbf{y}(t)-\textbf{y}(t-1))(\textbf{y}(t+1)-\textbf{y}(t))<0,$ (6)
$~{}\textbf{y}(t)>\textbf{y}(t-1)\},$
@@
$S_{v}(\textbf{y})=\{t\mid(\textbf{y}(t)-\textbf{y}(t-1))(\textbf{y}(t+1)-\textbf{y}(t))<0,$ (8)
$~{}\textbf{y}(t)<\textbf{y}(t-1)\}.$
@@
For the training of BBP-Net, we define $L_{BP}$ and $L_{ABP}$. $L_{BP}$ employs Huber loss [46] to calculate the loss of predicted SBP and DBP, and it is defined as the following equation.
@@
where $\hat{y}_{max}$ and $\hat{y}_{min}$ are the maximum and minimum values of $\hat{\textbf{y}}_{a}$. $L_{ABP}$ is defined as the L2 distance between the scaled acral rPPG signal $\hat{\textbf{y}}_{s}$ and its corresponding ABP signal $\textbf{y}_{ABP}$ as follows.
$L_{ABP}=\lVert\hat{\textbf{y}}_{s}-\textbf{y}_{ABP}\rVert_{2}.$ (17)
@@ IV Experimental results
Experiments were conducted on a hardware environment including Intel Core i9-10940X CPU, 64 GB DDR4 RAM, and NVIDIA Geforce RTX 3090 Ti. Pytorch was utilized to implement the proposed algorithm, and our code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG. In experiments, the temporal window was set to 150 samples which corresponds to 6 seconds. The learning rates of the Adam optimizer were set to 0.001 and 0.0001 for the MMSE-HR and V4V database, and the batch size was set to 8. To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE), root mean squared error (RMSE), and Pearson correlation coefficient r as evaluation measures.
IV-A Datasets
@@
We conducted experiments using the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) [16] and V4V (Vision for Vitals) database [17]. The datasets are sub-datasets derived from the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous ABP signals. Before collecting physiological data, each subject signed an informed consent form in accordance with the IRB approved protocol. As explained in Section III-B, pseudo PPG signals, heart rate, SBP, and DBP were obtained from ABP signals. In addition, the ABP measurement device used in these datasets is the Biopac NIBP100D, which can non-invasively measure continuous ABP signals by calibrating the finger PPG signals using cuff data. RGB video was recorded at the resolution of 1040$\times$1392 with the frame rate of 25 FPS, and ABP signals were collected at the sampling rate of 1000 Hz. The MMSE-HR database includes 102 video sequences from 17 male and 23 female subjects, and the average length of sequences is 30 seconds. The V4V database includes 1,358 data sequences from 179 subjects, and they are split into 724 for training, 276 for validation, and 358 for test samples. The average length of the sequences in the V4V database is 40 seconds.
During the acquisition of the MMSE database, subjects performed various tasks to arouse target emotions. The MMSE-HR dataset contains tasks designed to arouse emotions such as amusement, physical pain, anger, and disgust, while the V4V dataset additionally includes surprise, sadness, startle, skepticism, embarrassment, and fear. Specifically, tasks for surprise, startle, and embarrassment can induce large head motions, making the V4V dataset more challenging.
IV-B Experimental results on MMSE-HR database
@@
The performance of the DRP-Net for estimating heart rate is compared with previous methods on the MMSE-HR dataset. Following the previous work [29], experiments on the MMSE-HR dataset were conducted by using 5-fold cross-validation method for heart rate estimation. The folds were split independently between subjects to demonstrate generalizability of the proposed method. Table II presents the averaged performance over 5 folds for estimating heart rate from facial rPPG signals. The proposed method achieved MAE of 1.78, RMSE of 4.27, and r of 0.95, respectively, outperforming previous deep learning models with a significant margin. Previous methods proposed by Nowara et al. [27], Jaiswal & Meenpal [47], Ouzar et al. [13] infer rPPG signals based on small window sizes with low latency. However, insufficient temporal information leads to increase errors in estimating rPPG signals, resulting in higher heart rate estimation errors. Video transformer-based models proposed by Yu et al. [12] and Yu et al.[29] demonstrate significant improvements in cross-dataset tests. However, transformers require high-performance computing resources due to their large number of parameters and computational demands. As shown in Table II, our proposed model achieved better performance compared to the previous methods by using similar length of video sequences. In addition, the heart rate estimation errors from acral rPPG signals are 1.91, 4.74, and 0.93, respectively. DRP-Net estimates two rPPG signals with different phases and minimal errors for utilization in the blood pressure estimation stage.
TABLE II: Heart rate estimation results on the MMSE-HR database.
Method Window (s) MAE (BPM) RMSE r
POS [25] - 5.77 - 0.82
DeepPhys [11] - 4.72 8.68 0.82
Benefit of distraction [27] 2 2.27 4.90 0.94
EfficientPhys-C [48] - 3.48 7.21 0.86
CAN with synthetic data [49] 30 2.26 3.70 -
PhysFormer [12] 6.4 2.84 5.36 0.92
Spatiotemporal feature [47] 5 6.40 6.82 0.95
X-iPPGNet [13] 2 4.10 5.32 0.85
PhysFormer++ [29] 6.4 2.71 5.15 0.93
CIN-rPPG [50] 12 1.93 4.43 0.94
Dual-TL [51] 12 2.25 4.27 0.93
Ours 6 1.78 4.27 0.95
@@
Table III presents the results of estimating blood pressure on the MMSE-HR database. The proposed BBP-Net for blood pressure estimation outperformed previous methods, achieving MAE and RMSE of 10.19 and 13.01 for SBP and 7.09 and 8.86 for DBP, respectively. The algorithms proposed by Rong & Li [52] and Schrumpf et al. [53] extracted rPPG signals using a non-parametric approach and estimated blood pressure based on handcrafted features and deep learning models. These previous studies are similar to our approach in the aspect of extracting rPPG signals in an intermediate step. However, the main difference of our proposed method is to analyze phase-shifted rPPG signals extracted from DRP-Net. Chen et al. [36] proposed a blood pressure estimation model that utilizes two-dimensional spatiotemporal maps obtained from facial videos. However, this previous method cannot analyze pulse transit time of rPPG signals and shows insufficient performance for estimating SBP and DBP. Experimental results in Table III demonstrate that analyzing temporal discrepancy in phase-shifted rPPG signals is meaningful for improving the precision of blood pressure estimation.
TABLE III: Blood pressure estimation results on the MMSE-HR database.
Method Window (s) SBP DBP
@@
NCBP [52] - 17.52 22.43 12.13 15.23
NIBPP [53] 7 13.60 - 10.30 -
BPE-Net [36] 6 12.35 16.55 9.54 12.22
@@
Fig. 6 presents the Bland-Altman plots of predicted and reference blood pressure for the MMSE-HR database. In Fig. 6, solid line and dotted lines represent mean error and 95% limits of agreement, respectively. The results for SBP and DBP show positive errors at higher blood pressure and negative errors at lower blood pressure. These results indicate a bias in the estimated blood pressure and suggest that the generalization performance of the blood pressure estimation model is insufficient. To improve the generalization performance of the model, it is essential to collect more uniform dataset and broaden its distribution. Future work will focus on constructing datasets that include sufficient hypotensive and hypertensive data to enhance the generalizability of blood pressure estimation models.
@@
IV-C Experimental results on V4V database
This section presents experimental results for estimating heart rate on the V4V database. In Table IV, the performance of our proposed model is compared with previous methods. Facial rPPG signals estimated from the DRP-Net outperformed others in terms of MAE and r, with the values of 3.83 and 0.75, respectively. The performance comparison between PhysNet [28] and our DRP-Net indicates that utilizing wider receptive fields through atrous convolution is more advantageous for learning spatiotemporal features than the encoder-decoder structure based on three dimensional CNN. Moreover, compared to APNET [54], our facial rPPG signals showed lower error rates in terms of MAE and r. For the heart rate estimation results derived from acral rPPG signals, we achieved errors of 4.13, 10.14, and 0.73, respectively.
TABLE IV: Heart rate estimation results on the V4V database.
Method Window (s) MAE (BPM) RMSE r
DeepPhys [11] 30 10.20 13.25 0.45
PhysNet [28] - 13.15 19.23 0.75
APNET [54] - 4.89 7.68 0.74
Ours 6 3.83 9.59 0.75
Table V presents the results of estimating blood pressure on the V4V database. Our proposed algorithm achieved MAE and RMSE of 13.64 and 16.78 for estimating SBP and of 9.40 and 11.90 for estimating DBP, respectively. There is a scarcity of literature reporting on the performance of blood pressure estimation using the V4V database. We compared the performance of the proposed method to previous algorithms proposed by Schrumpf et al. [53] and Hamoud et al. [55]. As shown in Table V, our BBP-Net achieved lower MAE and RMSE in estimating both SBP and DBP, with significant margins.
TABLE V: Blood pressure estimation results on the V4V database.
Method Window (s) SBP DBP